OpenStudy (goldie1962):

If the value, to the nearest thousandth, of cos θ is -0.385, which of the following could be true about θ? A. 0 ≤ θ < π/6 B. π/6 ≤ θ < π/3 C. π/3 ≤ θ < π/2 D. π/2 ≤ θ < 2π/3 E. 2π/3 ≤ θ ≤ π My ACT Booklet gave me the answer D and an explanation, but I don't understand the explanation at all. Please help.

OpenStudy (campbell_st):

ok... so cos is negative... which quadrants are you in..?

OpenStudy (goldie1962):

Well in the fourth quadrant cos is positive, right?

OpenStudy (campbell_st):

well your question says \[\cos(\theta) = -0.385\] is that the correct question, or is it 2. \[\cos(\theta) = 0.385\]

OpenStudy (goldie1962):

Yes cos(θ) = -0.385

OpenStudy (campbell_st):

ok... so which quadrants is cos negative..? that is the key to the question. 4th and 1st quadrants cos is positive

OpenStudy (goldie1962):

Either second or third I think

OpenStudy (goldie1962):

Okay well I can eliminate A, B, C because they occur in the first quadrant. But doesn't E also fit the bill?

OpenStudy (campbell_st):

ok... so the next setp is find the angle \[\theta = \cos^{-1}(0.385)\] then 2nd quadrant angles are \[\pi - \theta or 180 - \theta... \] so you should be able to find the correct solution from there

OpenStudy (campbell_st):

so interms of angles measured in degrees D is between 90 and 120 E is between 120 and 180

OpenStudy (goldie1962):

Thank you! I got approx 113 degrees. Why is it π - θ, not 2π/3 - θ?

OpenStudy (loser66):



OpenStudy (loser66):

only D or E is the answer. It is D, not E, why? because the beginning of option E is 2pi/3